OpenStudy (anonymous):

In order to solve the literal equation C equals five ninths times F minus thirty-two for F, what would you do to the equation? Add 32 to both sides. Then, multiply both sides by nine fifths. Multiply both sides by nine fifths. Then, add 32 to both sides. Subtract 32 from both sides. Then, multiply both sides by nine fifths. Multiply both sides by nine fifths. Then, add 32 to both sides.

OpenStudy (whpalmer4):

\[C = \frac{5}{9}F-32\]If you are trying to solve for F, you want F alone on one side of the equals sign, and everything else on the other side. Generally you work backwards in the order of precedence: add and subtract, then multiply and divide, etc. Here before we can get rid of the 5/9 in front of F we must first get rid of the -32. How do we do that? Well, -32 + 32 = 0, so we add 32 to each side (must do the same thing to each side of the equation to preserve the equality). That gives us\[C+32 = \frac{5}{9}F-32+32\]or\[C+32 = \frac{5}{9}F\]Now we need to multiply or divide to isolate F.